I have really tried answering this problem, could you please help me out I would really appreciate it.
Question: If angle A and B are complementary angles,and angle A is 9 degrees less than twice angle B, find the measures of angles A and B.

Answer by nycsharkman(136)

As soon as you see the word complementary, you must know that the angles add up to 90 degrees.
angle A + angle B = 90
angle A = 9 less than TWICE angle B. We don't know what angle B is, right? I will say that angle B = x.
angle A = 2x - 9
We now have this equation:
2x - 9 + x = 90
Solve for x.
3x - 9 = 90
3x = 90 + 9
3x = 99
x = 99/3
x = 33
The measure of angle A = 2(33) - 9.
angle A = 66 - 9
angle A = 57 degrees
===================================
angle B = x and I just found that x = 33, right?
Then, angle B = 33 degrees.
How do you know that I am right?
When adding 33 + 57, I get 90 degrees ON BOTH SIDES of the equation.
Did you follow?
NOTE: If you see a similar question saying that two angles are supplementary, this means that both angles, when added, will produce 180 degrees. I give you this information now because most likely you will a question like that in your math course.